OpenStudy (vishweshshrimali5):

What does versin function mean .. ? The question is prove that...... 2 versin A + \(cos^2 A\) = 1 + \(versin^2 A\)

OpenStudy (vishweshshrimali5):

@Ishaan94 @mukushla @annas

OpenStudy (vishweshshrimali5):

I just want to know the meaning........ Then I would try the proof myself

OpenStudy (anonymous):

its 1 − cos(θ)

OpenStudy (anonymous):

yeah \(\text{versin} \ \theta=1-\cos \theta\)

OpenStudy (vishweshshrimali5):

Thanks a lot .......... everyone...........

OpenStudy (anonymous):

yw :)